OpenStudy (anonymous):

Mr.Daniels had a sum of money to invest. he invest 1/2 of the sum at 5%, 1/3 of the sum at 6% and the remainder at 3%. his total annual income was $300. what was the total sum that Mr.Daniels invested?

OpenStudy (anonymous):

solve \(.05\times \frac{1}{2}x+.06\times \frac{1}{3}x+.03\times \frac{1}{6}x=300\) for x